OpenStudy (anonymous):

Suppose n=(a_k)10^k+(a_k-1)10^(k-1)+....+(a_1)10+a_0 where 0<=a_j<=9 for all j. Prove n==a_0+a_1+...+a_k(mod 9)

OpenStudy (anonymous):

Suppose \[n=a _{k}10^{k}+a _{k-1}10^{k-1}+...+a _{1}10+a _{0}\] where \[0 \le a _{j} \le 9\] Prove the following: \[n =a _{0}+a _{1}+....+a _{k} \left(\mod9 \right)\]

OpenStudy (anonymous):

the only difference is that the equal sign is suppose to be the one with 3 bars instead of 2